a rectangle is three times as long as it is wide. If the perimeter is 104 feet, what are the length and width
Answer by checkley77(12844)

L=3W
2L+2W=104
2*3W+2W=104
6W+2W=104
8W=104
W=104/8
W=13 FT. ANS. FOR THE WIDTH.
L=3*13=39 FT ANS. FOR THE LENGTH.
PROOF:
2*39+2*13=104
78+26=104
104=104